# Artificial Intelligence HCI Human-Computer Interaction

### v2

@@ 6.3. Limitations & Future Work
Our sample consisted of participants with varying levels of experience in qualitative coding and different data types, which limited our ability to provide a unified, in-depth account of one specific QDA approach. Still, the diversity of reported QDA practices highlights the complexity of AI integration. Additionally, all participants were from HCI research, and many also specialized in security and privacy, potentially skewing our findings toward those more familiar with (emerging) technologies and computing. We similarly recognize that participants’ related motivational factors (Bandura, 2012), their differential attitudes (Ajzen et al., 2019), even their pre- or misconceptions (Gawronski and Strack, 2012; Hammer, 1996), and incidental knowledge (Ferguson and Bargh, 2024) concerning AI tools, can influence their responses. Nonetheless, our sample provided detailed insights into the potential for adaptive AI tools that can cater to both novice guidance and expert customization.
@@ Appendix C Examples of Participants’ Coding Workflow Drawings
Figure 15. Four examples of coding workflow drawings created by our participants during interviews. The drawings varied considerably in the level of detail they provided. Some included in-depth descriptions of specific actions during each step, while others focused more on visuals, providing details of the process only verbally.:

### v3

@@ 7.3. Limitations & Future Work
Our sample consisted of participants with varying levels of experience in qualitative coding and different data types, which limited our ability to provide a unified, in-depth account of one specific QDA approach. Still, the diversity of reported QDA practices highlights the complexity of AI integration. Additionally, our sample is slightly skewed towards early career researchers, and all participants were from HCI research, many also specialized in security and privacy, potentially skewing our findings toward those more familiar with and open to usage of (emerging) technologies and computing. Thus, they represent potential early adopters, making their insights and concerns essential. We did not collect data on prior exposure to AI or LLMs, but recognize that participants’ related motivational factors (Bandura, 2012), their differential attitudes (Ajzen et al., 2019), even their pre- or misconceptions (Gawronski and Strack, 2012; Hammer, 1996), and incidental knowledge (Ferguson and Bargh, 2024) concerning AI tools, can influence their responses. Nonetheless, our sample provided detailed insights into the potential for adaptive AI tools that can cater to both novice guidance and expert customization. Due to our sampling method choice (particularly relying on snowball sampling through our professional networks), our sample consisted of people working in a WEIRD context. This inherently limits the generalizability of our findings as well as how globally representative they are. Future work can and should interview researchers working in non-WEIRD contexts to bring in additional perspectives regarding how AI can and should be involved in QDA.
@@ Appendix C Examples of Participants’ Coding Workflow Drawings
Figure 14. Four examples of coding workflow drawings created by our participants during interviews. The drawings varied considerably in the level of detail they provided. Some included in-depth descriptions of specific actions during each step, while others focused more on visuals, providing details of the process only verbally.
This figure shows four different photos of workflow drawings created by participants during the interviews. One workflow is very complex, containing many substeps. The second workflow includes only three main steps (Interviews, followed by coding using tool support, and finally, the paper itself). The participant wrote short notes, such as ”intercoder agreement” and ”thematic analysis,” next to the arrows. The third workflow contains only icons and no text. It depicts a six-step process: 1) a piece of paper and a broom, 2) a piece of paper, a pen, and eyes, 3) two speech bubbles, 4) a mind map, 5) three people, and 6) a piece of paper labeled ”Results.” The last picture is simpler; it contains only a circle consisting of four steps: Collect data, meet and discuss, code, and iterate codebook. Below this, the participant created a small table indicating the columns their typical codebook would have.